Question:
4 (y-1) + 3 = 2y - 7
What I have done so far:
4y + 3 = 2y - 7 <-----the second step
I got this question from Mathpower 8

Answer by checkley71(8403)

4(y-1)+3=2y-7
4y-4+3=2y-7 (you forgot the -4 term)
4y-2y=-7+4-3
2y=-6
y=-6/2
y=-3 answer.